12. Sarah is exactly 12 years older than twice Jimmy's age. How many years from now will Sarah be only twice Jimmy's age?

Answer by lwsshak3(11628)

Sarah is exactly 12 years older than twice Jimmy's age. How many years from now will Sarah be only twice Jimmy's age?
***
At present:
x=Jimmy's age
2x+12=Sarah's age
let y=Years Sarah be only twice Jimmy's age.
2x+12+y=2(x+y)
2x+12+y=2x+2y
y=12
Years from now Sarah will be only twice Jimmy's age=12
Note: It doesn't matter what their present ages are.